Given the following experimental data, find the raPLEASE SHOW ALL OF YOUR WORK! THANK YOU.
Answer #1

The given reaction is NO(g) + NO2(g) + O2(g) ----> N2O5(g)

Let the rate law be r = k[NO]x[NO2]y[O2]z     ----(1)

Where

x = order of the reaction with respect to NO

y = order of the reaction with respect to NO2

z = order of the reaction with respect to O2

From the table of content ,

From (1) & (2) , On doubling the concentration of NO(2x0.10= 0.20M) kept the concentration of the NO2 & O2 constant,the rate of the reaction becomes doubles(2x2.1x10-2 = 4.2x10-2 ) so the value of x = 1

From (1) & (4) , On doubling the concentration of O2(2x0.10= 0.20M) kept the concentration of the NO2 & NO constant,the rate of the reaction becomes the same (1x2.1x10-2 = 2.1x10-2 ) so the value of z = 0

From (2) & (3) , On tripling the concentration of NO2 (3x0.10= 0.30M) & doubling the concentration of O2(2x0.1=0.20M)( it has no effect on rate from above) kept the concentration of the NO constant,the rate of the reaction becomes triples(3x4.2x10-2 = 1.26x10-1 ) so the value of y = 1

Therefore the rate law be r = k[NO]1[NO2]1[O2]0

   So the rate law is r = k[NO][NO2]     ----(2)

Plug the first values in the given tabular form we get

2.1x10-2 Ms-1 = k(0.10M ) x ( 0.10M)

      k = 2.1M-1 s-1

Therefore the rate constant is 2.1M-1 s-1